The binary operation on Z given by x
mathrmast (y=1+xy) not in the question later appered when i pasted)
is______ but not associative
a.Associative
b.Commutative
c.Distributive
d.Additive
1
Expert's answer
2020-05-26T19:21:56-0400

Let "y=1+xy" then

"xy=y-1"

The binary operation is not associative :

"(xy)z\neq x(yz)\\
(xy)z=(y-1)z=yz-z=z-1-z=-1\\
x(yz)=x(z-1)=xz-x=z-1-x"

The binary operation is not commutative: 

"xy\neq yx\\
xy=y-1\\
yx=x-1"

The binary operation is not distributive:

"(x+y)z\neq xz+yz\\
(x+y)z=z-1\\
xz+yz=xz+yz=z-1+z-1=2z-2\\
z(x+y)z\neq zx+zy\\
z(x+y)=x+y-1\\
zx+zy=x-1+y-1=x+y-2\\"

 The binary operation is additive

Answer d